OpenStudy (anonymous):

What are the intercepts of the line 5x+y=20 A. x - intercept 4 ; y - intercept 20 B. x - intercept 20 ; y - intercept -4 C. x - intercept -4 ; y - intercept -20 D. x - intercept -20 ; y - intercept -4

OpenStudy (zpupster):

first put it in a slope intercept form y=mx+b do this by subtracting 5x from both sides then let x = 0 y=? this is the y intercept

OpenStudy (anonymous):

can you go into more detail when you say both sides? @zpupster

OpenStudy (zpupster):

here 5x-5x+y=20-5x rearranging y=5x+20

OpenStudy (zpupster):

sorry imean y= -5x+20

OpenStudy (anonymous):

C? @zpupster

OpenStudy (zpupster):

plug in 0 for x

OpenStudy (zpupster):

plug in 0 for x